1. Copper shot (0.100 kg) is heated in boiling oil. After this, it is submerged in old water (0.300 kg at 20 ºC). If the water’s temperature turned 23.0 ºC, what was the temperature of the hot oil?

2. What is the new length of a bar made of pure gold at 75 ºC if it has a length 0.3 meters at 10 º C?

3.At what temperature will the Celsius and Fahrenheit scale have the same value?



1
Expert's answer
2022-02-09T06:27:58-0500

1.

"c_cm_c(t_f-t)+c_wm_w(t_f-t_w)=0,\\\space\\
t=\frac{c_wm_w(t_f-t_w)}{c_cm_c}+t_f=121°\text C."


2. "L=L_0(1+\alpha \Delta t)=0.3003\text{ m}."

3. The temperature is -40°C.